OpenStudy (anonymous_user):

Find the value of X: x^-2=1/16 A. 4 and -4 B. only 4 C. 16 and -16 D. only 16

OpenStudy (solomonzelman):

\[1/4^2=1/16\]

OpenStudy (anonymous):

\[\huge x^-2=\frac{1}{16} \rightarrow \frac{1}{x^2}=\frac{1}{16}\] \[\huge \rightarrow \frac{1}{x^2}=\frac{1}{4^2} \rightarrow \frac{1}{x}=\frac{1}{4}\] By invertendo x= 4

OpenStudy (solomonzelman):

you can use ^{ -2} as an exponent of -2.